OpenStudy (anonymous):

Solve the equation by graphing:-8x^3-13x^2+6x=0 a.no solution b.-2,.38 c.0,2,-.38 d.0,-2,.38

OpenStudy (anonymous):

1. divide the whole thing by x -8x^2-13x+6=0 -8x^2-16x+3x+6=0 -8x(x+2)+3(x+2)=0 (-8x+3)(x+2)=0 x1=3\8 x2=-2

OpenStudy (anonymous):

\[-8x^3-13x^2+6x = x (-8x^2 - 13x + 6)\] Then you have to use the quadratic formula to find the other two roots (one of the roots is x = 0)